# Embodied Instruction Following in Unknown Environments

### v1

@@ Abstract
Enabling embodied agents to complete complex human instructions from natural language is crucial to autonomous systems in household services. Conventional methods can only accomplish human instructions in the known environment where all interactive objects are provided to the embodied agent, and directly deploying the existing approaches for the unknown environment usually generates infeasible plans that manipulate non-existing objects. On the contrary, we propose an embodied instruction following (EIF) method for complex tasks in the unknown environment, where the agent efficiently explores the unknown environment to generate feasible plans with existing objects to accomplish abstract instructions. Specifically, we build a hierarchical embodied instruction following framework including the high-level task planner and the low-level exploration controller with multimodal large language models. We then construct a semantic representation map of the scene with dynamic region attention to demonstrate the known visual clues, where the goal of task planning and scene exploration is aligned for human instruction. For the task planner, we generate the feasible step-by-step plans for human goal accomplishment according to the task completion process and the known visual clues. For the exploration controller, the optimal navigation or object interaction policy is predicted based on the generated step-wise plans and the known visual clues. The experimental results demonstrate that our method can achieve 45.09% success rate in 204 complex human instructions such as making breakfast and tidying rooms in large house-level scenes.
1 Introduction
@@
To achieve this, end-to-end methods [24, 37, 32] directly generate the low-level actions from raw image input and natural language with the supervision of expert trajectories. To reduce the learning difficulties in the complex task, modular methods [7, 12, 22, 16], sequentially learn the instruction comprehension, state perception, spatial memory construction, high-level planning and low-level control to complete human goals. Since embodied agents are expected to complete more diverse and complex instructions, large language models (LLMs) are widely employed in EIF [17, 34, 8, 20, 27] due to their strong reasoning power and high generalization ability. However, existing methods can only generate plans in known environments where categories of all interactable objects in the scene are given to LLMs. Since the agent does not know the objects in the unknown environment, the generated plans are usually infeasible because of interacting with non-existing objects. Figure 1 (a) demonstrates an example of existing methods, where the agent is unaware that no bottles exist in the unknown environment. Interacting with the non-existent bottles based on the infeasible plan fails to accomplish the human goals of water serving.
@@
In this paper, we propose an EIF method for complex tasks in the unknown environment. Different from conventional methods that assumes knowing interactable objects in advance, our method navigates the unknown environment to efficiently discover objects that are relevant to the complex human requirements. Therefore, the embodied agent can generate feasible task plans in realistic indoor scenes where the locations and existence of objects are frequently changing. Figure 1 (b) also demonstrates the same example of water serving implemented by our method, and our agent efficiently discovers the mug and uses it as the receptacle of water because no bottles exist in the scene. We first construct a hierarchical EIF framework including the high-level task planner and the low-level exploration controller with multi-modal LLMs, which are finetuned by the large-scale generated trajectories of the complex EIF tasks. We then design a scene-level semantic representation map to depict the visual clues in the known area, through which the goals of the task planner and the exploration controller can be aligned to feasibly complete human instructions.
@@
More specifically, the goal of the task planner is to generate feasible plans for human instruction including navigation and manipulation in natural language. The task planner predicts the next step based on the semantic representation map and the task completion process. The exploration controller aims at discovering task-related objects with low action cost, which selects the optimal navigation policy from all navigable borders or object interaction policy according to the semantic representation map and the generated step-wise plans. For the scene-level semantic feature map, we project the CLIP features of collected RGB images during exploration to the top-down map with dynamic region attention, which preserves the task-relevant visual information in the map without redundancy. The experimental results in ProcTHOR and AI2THOR simulation environments show that our method can achieve 45.09% success rate in 204 complex human instructions in large house-level scenes.
2 Related Works
Embodied Instruction Following: The EIF task requires the robot to follow human instructions represented by natural language in the interactive environment. Recent research can be divided into two categories: end-to-end approaches and modular methods. For the former, agents utilize RGB-D image sequences to represent the scene and generate low-level interactive actions directly through image sequences and language embedding. Singh et al. [30] mitigated semantic confusion by providing global and pixel-level semantic cues for action prediction and object localization, and Nguyen et al. [23] exploited the relationship between instructions and visual clues. However, the end-to-end approaches require large-scale images and annotated low-level action data, resulting in low performance on unseen scenes. The modular approaches contain high-level planning, low-level control and spatial memory construction modules with enhanced unreliability. Min et al. [19] significantly improved task execution success via employing explicit spatial memory and semantic search strategies for agent guidance, and Kim et al. [15] enhanced the efficiency of the agent by focusing on task-relevant objects with the usage of contextual planning and memory frameworks. However, existing EIF methods are designed for known environments where visual clues of the whole scene can be easily acquired by looking around, and fail to discover required visual clues in unknown environments for feasible action generation.
@@ 2 Related Works
Scene Representation for Visual-language Navigation: Visual-language navigation requires agents to explore unknown environments to locate target objects and follow natural language instructions. The primary challenge lies in efficiently representing expansive unknown scenes for generating navigation policies. Existing scene representations consist of three categories: 2D semantic maps, 3D geometric maps and scene graphs. Early works [3, 2] constructed the 2D semantic maps by projecting visual clues in the top-down view, which are leveraged for navigation frontier selection for target finding. PONI [25] proposed a scoring network for all potential frontiers of unseen regions 2D semantic maps, and L3MVN [35] determined the semantic relevance of the objects around each frontier to the target by BERT [6]. To embed the geometric information, 3D geometric maps are investigated by fusing the structure and semantic information. LERF [14] and ConceptFusion [13] integrated fine-grained alignment of semantic features with 3D maps in SLAM, multi-view fusion, and NeRF [18] for multiple downstream tasks. To reduce the storage overhead, scene graphs [9, 11] are proposed to represent objects or concepts as nodes and spatial relations as edges to represent the scene topology efficiently. SayPlan [26] enabled agents to focus on task-relevant nodes by integrating subgraph folding and replanning mechanisms. Inspired by the above approaches, we construct semantic feature maps to empower embodied agents to explore unknown environments, where task-relevant information can be acquired for action generation with low exploration cost.
3 Problem Statement
@@
Given the human instruction $I$ in natural language, the robot should generate a sequence of action primitives including (PickUp, Place, Open, Close, ToggleOn, ToggleOff, Slice) to complete the instruction. The agent can only acquire the scene information for instruction following via an RGB-D camera mounted on the agent, through which the agents build a semantic map $S$ to generate the feasible interaction. In realistic deployment, the embodied agent usually work in unknown environments, where the location and existence of objects in the house-level scene are not known. Therefore, we add an additional action primitive (Navigation) to enable the agent to explore the scene for visual information collection.
@@
4 Approach
In this section, we first introduce the overall pipeline of our EIF method designed for unknown environments, and then we describe the details of the high-level planner and the low-level controller. Moreover, we elaborate the construction of the online semantic feature maps that ground the planner and the controller to the physical scene. Finally, we demonstrate the model training and the inference of our framework in practical deployment.
4.1 Overall Pipeline
@@
In realistic deployment scenarios of household robots, the physical world is usually unknown for the agent because the existence and locations frequently change due to human activity. Therefore, the agents are required to construct the online scene feature map according to the real-time visual perception during the robot navigation, through which the agent generates feasible step-by-step plans to achieve the human goal and the efficient exploration trajectories for the unknown scene including navigation and object interaction to complete each step in the plan. Figure 2 demonstrates the overall pipeline of our agent. The scene feature map represents the visual clues of the scene in the top-down view based on the collected RGB-D images during exploration, where the pre-trained features of regions with higher relevance to the instruction are assigned with higher importance for feature map construction. The high-level planner generates the plans for the next step with natural language based on the task completion process and the semantic feature map, and the low-level controller predicts the templated action primitives, location and target objects for executable navigation or manipulation based on the scene feature map and the plan for the next step.
4.2 Hierarchical Embodied Agents for EIF in Unknown Environments
We decompose EIF in unknown environments into two sub-tasks including the high-level planning and the low-level exploration. The generated high-level plans are leveraged as guidance for the agent to select the most relevant regions for exploration, and the predicted low-level actions update the semantic feature maps to provide visual clues for feasible plan generation. Both the planner and the explorer are implemented by a finetuned LLaVA model.
@@ 4.2 Hierarchical Embodied Agents for EIF in Unknown Environments
High-level planner: The planner generates the plan for the next step in natural language, which considers the textual information including the human instruction and the completed steps and the visual clues represented by the semantic feature maps. The forward pass of the high-level planner $HP$ can be represented as follows:
$p_{i}=HP(I,\{p_{k}\}_{k=1}^{i-1};S_{i-1})$ (1)
where $S_{i}$ means the semantic feature maps updated in the $i_{th}$ step and we leverage a LLaVA model whose visual encoder is the ViT-L/14 architecture for the high-level planner.
@@
Low-level controller: The low-level controller predicts the specific actions including the action primitives, locations, and target objects according to the generated high-level plans and the semantic feature maps, which explores the unknown scene and completes the step-wise plan. The forward pass of the low-level controller $LC$ can be represented as follows:
$\{a_{j}^{i},l_{j}^{i},o_{j}^{i}\}=LC(p_{i},\{f_{m}^{i}\}_{m};\{s_{m}^{i}\}_{m})$ (2)
@@
where $l_{j}^{i}$ and $o_{j}^{i}$ are the predicted location and target objects for the $j_{th}$ actions in the $i_{th}$ step of the high-level plan. Meanwhile, $f_{m}^{i}$ means the textual features of the $m_{th}$ segment of the frontier between known and unknown regions for $S_{i}$, where $m$ represents the number of frontier segments in the entire $S_{i}$. The textual features are demonstrated by the coordinate of the middle point for the frontier segment. $s_{m}^{i}$ denotes the semantic features of the $m_{th}$ frontier segments, which is demonstrated by the semantic feature map patches containing the corresponding frontiers. The low-level controller not only explores the unknown scene with navigation and object interaction but also completes the step-wise plans by manipulating the target object (e.g. pick up the tomato). For action primitives except for navigate, the predicted actions are implemented on the target objects. For navigate, the robot just moves to the predicted locations without object interaction.
4.3 Online Semantic Feature Maps
@@
The high-level planner and the low-level controller should be aligned so that they can generate feasible plans and exploratory actions to achieve human instructions in the unknown environment. The semantic feature maps can be leveraged for alignment since they provide visual clues of the scene for both the high-level planner and the low-level controller. In realistic deployment scenarios of household robots, the existence and locations frequently change due to human activity. Therefore, we propose an online semantic feature map that is dynamically updated during the exploration of the unknown scene for each human instruction.
Semantic feature maps mean the visual representation for each pixel in the top-down view of the entire house. Compared with conventional semantic maps, our semantic feature maps can represent implicit relationships between objects in the scene instead of only categories, which provides crucial information for effective exploration policy generation. For EIF in unknown environments, the visual information collected in the $i_{th}$ timestep contains the RGB image $C_{i}$ and the depth image $D_{i}$. To enable the semantic feature maps to acquire high generalization ability in diverse human instructions, we leverage CLIP to extract the pixel-wise visual features $\mathbf{f}_{xy}^{i}$ at time $i$ for the pixel in $x_{th}$ row and $y_{th}$ column of $C_{i}$ via following ConceptFusion [13], and fuse the feature of the entire image and that of the instance mask for the pixel as the pixel-wise visual features $\mathbf{f}_{xy}^{i}$ in the same way. The visual features contribute to the projected location in the scene feature map in the top-down view, which can be depicted as follows:
$\mathbf{F}_{uv}^{i}=\sum_{x,y}\mathbf{f}_{xy}^{i}\cdot\mathbb{I}(\mathcal{P}_{D_{i}}(x,y)\in\mathcal{S}(u,v))$ (3)
@@
where $\mathbf{F}_{uv}^{i}$ means the contribution to the element in the $u_{th}$ row and $v_{th}$ column of the semantic feature map from the visual information collected in time $i$, and $\mathcal{P}_{D_{i}}(x,y)$ demonstrates the projected coordinates in the top-down view for of the pixel $(x,y)$ based on the depth image $D_{i}$. $\mathcal{S}(u,v)$ means the pixel in the $u_{th}$ row and $v_{th}$ column in the semantic feature map, and The indicator function $\mathbb{I}(\cdot)$ equals one for true and zero otherwise.
@@
The semantic feature map is updated at each time step during the exploration process, where the agent sees new visual information for recording. Since the house for embodied instruction following in realistic world is usually very large, regarding all images with equal importance in semantic feature map construction leads to significant information redundancy. Meanwhile, different visual clues usually make various contribution to the given human instruction. Therefore, we should assign large importance to relevant visual clues when updating the semantic feature maps, so that sufficient visual information can be represented without redundancy for high-level planning and low-level exploration. The task relevance can be acquired as follows. The high-level planner is also required to generate the demanded objects $\{O_{k}\}_{k}$ for the predicted corresponding step-wise plan, which are leveraged to construct three prompts including (a) the image contains $\{O_{k}\}_{k}$, (b) the image does not contain $\{O_{k}\}_{k}$ and (c) the image contains nothing. We then leverage a pre-trained LongCLIP [36] to predict the similarity score between the image and all prompts. Finally, the online semantic feature map is updated with dynamic region attention:
$\mathbf{S}_{uv}^{i}=(1-w_{i})\mathbf{S}_{uv}^{i-1}+w_{i}\mathbf{F}_{uv}^{i}$ (4)
@@
where $\mathbf{S}_{uv}^{i}$ means the features in the $i_{th}$ row and $j_{th}$ column of the semantic feature maps at time $i$. The normalized weight $w_{i}=\hat{w_{i}}/\frac{1}{i}\sum_{k=1}^{i}\hat{w}_{k}$ represents the importance of the current semantic features compared with known visual clues, where $w_{k}$ is the original similarity score between the image and the prompt in the $k_{th}$ time step. The online semantic feature maps contain rich visual information, and the most relevant regions can be explored via navigating the optimal border and interacting with related objects to achieve human goals with minimized action cost.
4.4 Training and Inference
Training: The training samples for the high-level planner consist of human instruction, current completed plans, current semantic feature maps and the groundtruth plan for the next step, and those for the low-level controller include plan for the next step, textual and semantic features for current border segments and the groundtruth action sequences representing primitives, location and targets. The details of input and output are provided in Appendix E. We leverage GPT-4 and the ProcTHOR simulator to generate the large-scale dataset to train the LLaVA-based high-level planner. We annotate several seed instructions and leverage GPT-4 to generate more instructions and corresponding plans based on the object list for each scene in the ProcTHOR, where samples with logical errors are filtered with PDDL parameters [28]. We then implement the generated plans in ProcTHOR and collect the navigation trajectories, RGB-D images, object locations and robot poses as the training data. Finally, the generated samples are parsed into high-level planning samples and low-level action data. We follow the supervised fine-tuning paradigm in LLM for training the LLaVA model in high-level planner and low-level controller, where we mask out $p_{i}$ and $\{a_{j}^{i}\}_{j=1}^{\tau_{i}}$ in the $i_{th}$ step. The CE loss leveraged in the training process is represented by:
where $\boldsymbol{X}_{V}$ denotes scene feature maps and $\boldsymbol{X}_{T}$ means input text prompt tokens. $\boldsymbol{R}_{<m}$ represents the output text tokens before the $m_{th}$ token $R_{m}$ and $M$ arenumber of output tokens. In this way, the pre-trained multimodal LLMs can be grounded to high-level planning and low-level control tasks in realistic scenes, where executable plans and actions are generated based on the scene representation.
Inference: The high-level planner generates the planning for the next step based on the current RGB-D image and scene information represented by the semantic map, and the low-level controller predicts the action primitives, target object and interaction position based on the generated step-wise plan. The semantic feature maps are updated when implementing the low-level action sequences. The high-level planner will generate the plans for the next step only when the current low-level action is successfully achieved. The detailed process is illustrated in Appendix B.
@@ 4.4 Training and Inference
Table 1: Comparison with different EIF methods across different instructions in the ProcTHOR simulator, where LLM-P represents the LLM-P without performing re-planning.
Method Small-scale Large-scale

### v2

@@ Abstract
Enabling embodied agents to complete complex human instructions from natural language is crucial to autonomous systems in household services. Conventional methods can only accomplish human instructions in the known environment where all interactive objects are provided to the embodied agent, and directly deploying the existing approaches for the unknown environment usually generates infeasible plans that manipulate non-existing objects. On the contrary, we propose an embodied instruction following (EIF) method for complex tasks in the unknown environment, where the agent efficiently explores the unknown environment to generate feasible plans with existing objects to accomplish abstract instructions. Specifically, we build a hierarchical embodied instruction following framework including the high-level task planner and the low-level exploration controller with multimodal large language models. We then construct a semantic representation map of the scene with dynamic region attention to demonstrate the known visual clues, where the goal of task planning and scene exploration is aligned for human instruction. For the task planner, we generate the feasible step-by-step plans for human goal accomplishment according to the task completion process and the known visual clues. For the exploration controller, the optimal navigation or object interaction policy is predicted based on the generated step-wise plans and the known visual clues. The experimental results demonstrate that our method can achieve 45.09% success rate in 204 complex human instructions such as making breakfast and tidying rooms in large house-level scenes. Code and supplementary are available at https://gary3410.github.io/eif_unknown/.
I INTRODUCTION
@@
To achieve this, end-to-end methods [12] directly generate the low-level actions from raw image input and natural language with the supervision of expert trajectories. To reduce the learning difficulties in the complex task, modular methods [5], sequentially learn the instruction comprehension, state perception, spatial memory construction, high-level planning and low-level control to complete human goals. Since embodied agents are expected to complete more diverse and complex instructions, large language models (LLMs) are widely employed in EIF [17] due to their strong reasoning power and high generalization ability. However, existing methods can only generate plans in known environments where categories of all interactable objects in the scene are given to LLMs. Since the agent does not know the objects in the unknown environment, the generated plans are usually infeasible because of interacting with non-existing objects. Fig. 1 (a) demonstrates an example of existing methods, where the agent is unaware that no bottles exist in the unknown environment. Interacting with the non-existent bottles based on the infeasible plan fails to accomplish the human goals of water serving.
@@
In this paper, we propose an EIF method for complex tasks in the unknown environment. Different from conventional methods that assume knowing interactable objects in advance, our method navigates the unknown environment to efficiently discover objects that are relevant to the complex human requirements. Therefore, the embodied agent can generate feasible task plans in realistic indoor scenes where the locations and existence of objects are frequently changing. Fig. 1 (b) also demonstrates the same example of water serving implemented by our method, and our agent efficiently discovers the mug and uses it as the receptacle of water because no bottles exist in the scene. We first construct a hierarchical EIF framework including the high-level task planner and the low-level exploration controller with multi-modal LLMs, which are finetuned by the large-scale generated trajectories of the complex EIF tasks. We then design a scene-level semantic representation map to depict the visual clues in the known area, through which the goals of the task planner and the exploration controller can be aligned to feasibly complete human instructions.
@@
More specifically, the goal of the task planner is to generate feasible plans for human instruction including navigation and manipulation in natural language. The task planner predicts the next step based on the semantic representation map and the task completion process. The exploration controller aims at discovering task-related objects with low action cost, which selects the optimal navigation policy from all navigable borders or object interaction policy according to the semantic representation map and the generated step-wise plans. For the scene-level semantic feature map, we project the CLIP features of collected RGB images during exploration to the top-down map with dynamic region attention, which preserves the task-relevant visual information in the map without redundancy. The experimental results in ProcTHOR [3] simulation environment show that our method can achieve 45.09% success rate in 204 complex human instructions in large house-level scenes.
II Related Works
Embodied Instruction Following: The EIF task requires the robot to follow human instructions represented by natural language in the interactive environment. A key challenge for the EIF task is generating interaction goals and actions grounded in the deployment environment according to the instructions. Prior works (e.g., LACMA [18], E.T. [12], M-TRACK [15]) have explored end-to-end transformer architecture to generate grounded low-level interaction actions based on the current environment perception, modular approaches (e.g., HLSM [2], FILM [10], LLM-Planner [16]) propose enhancing the generalization of unseen scenes with hierarchical planners. However, prior arts have focused on single-room environments, which are designed for known environments where visual clues of the whole scene can be easily acquired by looking around. The low scalability of the scene scale limits their ability to discover required visual clues in unknown environments for feasible action generation.
@@ II Related Works
Scene Representation for Visual-language Navigation: Existing scene representations consist of three categories: 2D semantic maps, 3D geometric maps and scene graphs. Early works [1] constructed the 2D semantic maps by projecting visual clues in the top-down view, which are leveraged for navigation frontier selection for target finding. L3MVN [21] determined the semantic relevance of the objects around each frontier to the target by BERT [4]. To embed the geometric information, 3D geometric maps are investigated by fusing the structure and semantic information. ConceptFusion [7] integrated fine-grained alignment of semantic features with 3D maps in SLAM, multi-view fusion, and NeRF [9] for multiple downstream tasks. To reduce the storage overhead, scene graphs [13, 20] are proposed to represent objects or concepts as nodes and spatial relations as edges to represent the scene topology efficiently. Inspired by the above approaches, we construct semantic feature maps to empower embodied agents to explore unknown environments.
III Problem Statement
@@
Given the human instruction $I$ in natural language, the robot should generate a sequence of action primitives including (PickUp, Place, Open, Close, ToggleOn, ToggleOff, Slice) to complete the instruction. The agent can only acquire the scene information for instruction following via an RGB-D camera mounted on the agent, through which the agents build a semantic map $S$ to generate the feasible interaction. In realistic deployment, the embodied agent usually works in unknown environments, where the location and existence of objects in the house-level scene are not known. Therefore, we add an additional action primitive (Navigation) to enable the agent to explore the scene for visual information collection.
@@
IV Approach
IV-A Overall Pipeline
@@
In realistic deployment scenarios of household robots, the physical world is usually unknown for the agent because the existence and locations frequently change due to human activity. Therefore, the agents are required to construct the online scene feature map according to the real-time visual perception during the robot navigation, through which the agent generates feasible step-by-step plans to achieve the human goal and the efficient exploration trajectories for the unknown scene including navigation and object interaction to complete each step in the plan. Fig. 2 demonstrates the overall pipeline of our agent. The scene feature map represents the visual clues of the scene in the top-down view based on the collected RGB-D images during exploration, where the pre-trained features of regions with higher relevance to the instruction are assigned higher importance for feature map construction. The high-level planner generates the plans for the next step with natural language based on the task completion process and the semantic feature map, and the low-level controller predicts the templated action primitives, location and target objects for executable navigation or manipulation based on the scene feature map and the plan for the next step.
IV-B Hierarchical Agents for EIF in Unknown Environments
@@
High-level planner: Formally, the high-level planner $P^{h}$ decomposes human instructions into step by step, generating the plan for the next step in natural language, which considers textual information including human instruction and completed steps and the visual clues represented by semantic feature maps.
$p_{i}=P^{h}(I,\{p_{k}\}_{k=1}^{i-1};S_{i-1})$ (1)
where $S_{i}$ means the semantic feature maps updated in the $i_{th}$ step.
@@
Low-level controller: The low-level controller predicts $P^{l}$ the specific actions, including action primitives, locations, and target objects according to the high-level plans generated and the semantic feature maps, which explore the unknown scene and complete the stepwise plan.
$\{a_{j}^{i},l_{j}^{i},o_{j}^{i}\}=P^{l}(p_{i},\{f_{m}^{i}\}_{m};\{s_{m}^{i}\}_{m})$ (2)
@@
where $l_{j}^{i}$ and $o_{j}^{i}$ are the predicted location and target objects for the $j_{th}$ actions in the $i_{th}$ step of the high-level plan. Meanwhile, $f_{m}^{i}$ means the textual features of the $m_{th}$ segment of the frontier between known and unknown regions for $S_{i}$, where $m$ represents the number of frontier segments in the entire $S_{i}$. The textual features are demonstrated by the coordinates of the middle point for the frontier segment. $s_{m}^{i}$ denotes the semantic features of the $m_{th}$ frontier segments, which is demonstrated by the semantic feature map patches containing the corresponding frontiers. The low-level controller not only explores the unknown scene with navigation and object interaction but also completes the step-wise plans by manipulating the target object (e.g. pick up the tomato). For action primitives except for navigate, the predicted actions are implemented on the target objects. For navigate, the robot just moves to the predicted locations without object interaction.
Both high-level $P^{h}$ and low-level $P^{l}$ are composed with VLMs, which leverage scene semantic information to decompose instructions into step-by-step plans and ground them into executable low-level actions.
IV-C Online Semantic Feature Maps
@@
The high-level planner and the low-level controller should be aligned so that they can generate feasible plans and exploratory actions to achieve human instructions in an unknown environment. In realistic deployment scenarios of household robots, the existence and locations frequently change due to human activity. Therefore, we propose an online semantic feature map that is dynamically updated during the exploration of the unknown scene for each human instruction.
Feature extraction: To enable semantic feature maps to acquire high generalization ability in various human instructions, we use vision foundation model to extract pixelwise visual features $\mathbf{f}_{xy}^{i}$ at time $i$ for the pixel in $x_{th}$ row and $y_{th}$ column of the RGB image $C_{i}$ by combining the feature of the entire image and that of the instance mask containing the corresponding pixel. The visual features contribute to the projected location in the scene feature map via the depth image $D_{i}$.
$\mathbf{F}_{uv}^{i}=\sum_{x,y}\mathbf{f}_{xy}^{i}\cdot\mathbb{I}(\mathcal{P}((x,y),D_{i})\in\mathcal{S}(u,v))$ (3)
@@
where $\mathbf{F}_{uv}^{i}$ means the contribution to the element in the $u_{th}$ row and $v_{th}$ column of the semantic feature map from the visual information collected in time $i$, and $\mathcal{P}$ demonstrates the projected function. $\mathcal{S}(u,v)$ means the pixel in the $u_{th}$ row and $v_{th}$ column in the semantic feature map, and the indicator function $\mathbb{I}(\cdot)$ equals one for true and zero otherwise.
Frontier representation: We generate frontier masks that distinguish between known and unknown regions based on the occupancy map for efficient exploration. Through connected component analysis, we obtain the mask of each frontier instance. We further remove frontiers with areas smaller than the threshold (150 pixels) to reduce redundant exploration. We sample 32 visual embeddings as frontier tokens according to the frontier instance mask on the corresponding region of the feature map, while utilizing the coordinates of their centroids for the frontier text description. The specific representation is illustrated in Fig. 3 (a).
@@
Dynamic region attention: Since the house for embodied instruction following in realistic world is usually very large, regarding all images with equal importance in semantic feature map construction leads to significant information redundancy. Meanwhile, different visual clues usually make various contribution to the given human instruction. Therefore, we should assign large importance to relevant visual clues when updating the semantic feature maps with task demanded objects $L_{dec}$, so that sufficient visual information can be represented without redundancy for high-level planning and low-level exploration. Specifically, we further expand $L_{dec}$ into $\overline{L_{dec}}$ and $L_{none}$ to match the input requirements of image and text alignment models such as CLIP. $\overline{L_{dec}}$ and $L_{none}$ describe the image as not containing the target objects and not containing the objects, respectively. The components as illustrated in Fig. 3 (b), and consider the score of $L_{dec}$ as the attention score $c_{i}$. Finally, the online semantic feature map is updated with dynamic region attention:
$\mathbf{S}_{uv}^{i}=(1-w_{i})\mathbf{S}_{uv}^{i-1}+w_{i}\mathbf{F}_{uv}^{i},\quad w_{i}=c_{i}/\frac{1}{i}\sum_{k=1}^{i}c_{k}$ (4)
@@
where $\mathbf{S}_{uv}^{i}$ means the features in the $i_{th}$ row and $j_{th}$ column of the semantic feature maps at time $i$. The normalized weight $w_{i}$ represents the importance of the current semantic features compared with known visual clues, where $c_{k}$ is the original similarity score between the image and the prompt in the $k_{th}$ time step.
@@
TABLE I: Comparison with different EIF methods across different instructions in the ProcTHOR simulator, where LLM-P represents the LLM-P without performing re-planning.
Method Normal-scale Large-scale